Given f(x)=xsquare +2x and g(x)=x square +5x-3,find the value fox such that f(x)=g(x).



Answer by Cromlix(4381)

x^2 + 2x = x^2 + 5x - 3
Combine:
x^2 - x^2 + 2x - 5x += -3
-3x = - 3
x = 1
Hope this helps.
:-)